OpenStudy (anonymous):

\[\int\frac{1}{\ln x}\,dx\]

OpenStudy (anonymous):

Now, I'm pretty sure\[u=\ln x\]\[1=\frac{dx}{du}\frac{1}{x}\]\[dx=x\,du=e^{\ln x}\,du=e^u\,du\]\[\int\frac{e^u}{u}\,du\]And I could swear that this can be expressed as some disgusting power series variant.

OpenStudy (anonymous):

\[\int\left(\frac{1}{u}e^u\right)\,du=\int\left(\frac{1}{u}\sum_{n=1}^\infty u^n\right)\,du\]

OpenStudy (anonymous):

@Zarkon Check this out please. :3

OpenStudy (anonymous):

\[\int\sum_{n=1}^\infty u^{n-1}\,du\]

OpenStudy (anonymous):

Whoops, forgot something!

OpenStudy (zarkon):

there is no antiderivative in terms of elementary functions

OpenStudy (anonymous):

Expressing the answer in terms of a sum would be fine.

OpenStudy (anonymous):

\[\int\frac{1}{u}e^u\,du=\int\frac{1}{u}\sum_{n=1}^\infty\frac{u^n}{n!}\,du=\int\sum_{n=1}^\infty\frac{u^{n-1}}{n!}\,du=\sum_{n=1}^\infty\frac{u^n}{n\cdot n!}\]

OpenStudy (anonymous):

@FoolForMath :D

OpenStudy (anonymous):

\[\sum_{n=1}^\infty\frac{u^n}{n\cdot n!}=\sum_{n=1}^\infty\frac{\ln(x)^n}{n\cdot n!}\]

OpenStudy (anonymous):

Can it be further simplified?

OpenStudy (anonymous):

A better question: what does\[\sum_{n=1}^\infty\frac{\ln(x)^n}{n\cdot n!}\]converge to?

OpenStudy (anonymous):

By the ratio test it converges.

OpenStudy (zarkon):

you should include your constant of integration along with the radious of convergence.

OpenStudy (anonymous):

just a correction :\[\int \frac{1}{u} \sum_{n=0}^{\infty} \frac{u^n}{n!} du=\int (\frac{1}{u}+\sum_{n=1}^{\infty} \frac{u^{n-1}}{n!})\ du=\ln u+\sum_{n=1}^{\infty} \frac{u^{n}}{n.n!}+C\]